OpenStudy (anonymous):

What can we afirm about the roots of 2x³ - x² - 2x + 1 = 0 ? a) only one of them is irrational b) they belong to [-1,1] c) two of them are negative d) none of them is real Can someone help me? I divided it by x - 1 and got x = -1 and x = 1/2. Is it right to say they belong to [-1,1]? Because option 'a' also seems right. Help!

sam (.sam.):

2x³ - x² - 2x + 1 = 0 has 3 roots x=1 x=-1 x=1/2

OpenStudy (anonymous):

Still, both a and b seem right.

sam (.sam.):

When x=-1 y=0 , it cant be b)

sam (.sam.):

a

OpenStudy (anonymous):

Yep, it can be b as well. why do you think its a?

sam (.sam.):

When x=-1 y=0 , it cant be b

sam (.sam.):

For [-1,1]

OpenStudy (anonymous):

I know that. Why did you post 'a'?

Directrix (directrix):

b) they belong to [-1,1] No roots for this polynomial are irrational. Do you know the Rational Roots Theorem, Synthetic Division? I don't know the divisor you used for your division or the dividend but if you provide that information or perhaps use the "Draw" button to write your division, I will try to help you.

OpenStudy (anonymous):

After finding 2x² + x - 1 I did this: x = - 1 + sqrt9 / 4 x = -1 - sqrt9 /4 Then I got x = -1 and x = 1/2

OpenStudy (anonymous):

1/2 is irrational, isn't it?

OpenStudy (anonymous):

oops haha

Directrix (directrix):

1/2 is rational. A rational number is one that can be written as the ratio of 2 integers. The first 5 letters of ratio nal are a reminder.

OpenStudy (anonymous):

Yep, sorry.

Directrix (directrix):

What did you divide by to get 2x² + x - 1? I'm trying to tune in to your thinking. I just saw this thread.

OpenStudy (anonymous):

I divided 2x³ - x² - 2x + 1 by x - 1

OpenStudy (anonymous):

remainder's 0 so I guess it's right

Directrix (directrix):

Okay, I got the same quotient. 2x^2 + x -1 should factor.

Directrix (directrix):

(2x - 1) (x + 1) = 0 2x - 1 = 0 or x + 1 =0 x = 1/2 or x = -1

OpenStudy (anonymous):

Thanks. I can see it clearly now, it's b indeed.

Directrix (directrix):

You understand the [-1,1] notation. Yes?

OpenStudy (anonymous):

Yep. Thx a lot

Directrix (directrix):

Great.